# State Permutation Control in Non-Hermitian Multiqubit Systems with Suppressed Non-Adiabatic Transitions

### v1

@@ III.2 Interacting qubits
When the qubit interactions in Eq. (1) are switched on, the NHH spectrum takes the form of a stack of Riemann sheet pairs, as illustrated in Fig. 2, with the spacing between sheets determined by a function of the $ZZ$ coupling strength $J_{mn}$. Hence, for interacting qubits, the EP of a single qubit extends to an exceptional line (EL), as depicted in Fig. 2. This EL is orthogonal to the $(x,y)$ plane and intersects it at the point $(x=0,y=1)$.
To determine whether a non-Hermitian multiqubit system enables a multistate switch, one can examine a topological invariant, associated with EPs, known as chirality [13]. Namely, when the chirality is non-zero, the eigenstates can be swapped while encircling an EP. In contrast, when the chirality is zero, winding around the EP does not induce a state flip; instead, each eigenstate returns to itself after completing a full encircling cycle [13].
@@
One can readily verify that, in general, for arbitrary values of magnetic fields $\vec{\mathbb{B}}_{||}^{i}$ and coupling strengths $J_{mn}$, the energy $E_{k}$, ($k=1,\dots,2N$), in the vicinity of each branch cut can be expanded as follows
$E_{k}=\begin{cases}\sum\limits_{j}u_{j}^{k}\alpha^{j},\quad&\text{when $N$ is odd},\\ \sum\limits_{j}u_{j}^{k}\alpha^{2j},\quad&\text{when $N$ is even},\end{cases}$ (10)
with $u_{j}^{k}\in{\mathbb{R}}$. From Eq. (10) one concludes that a non-zero chirality of the EPs and state transfer, associated with it, are always realizable in a system with an odd number of qubits.
We note that a non-zero chirality, and therefore the symmetric state transfer, for an even number of qubits can also be realized by making appropriate modifications to the NHH in Eq. (1). For example, in a two-qubit system, this can be achieved by applying a non-Hermitian synthetic magnetic field to one qubit, while the second qubit is subjected to a real-valued field.
@@ IV Implementing a state permutation group in a non-hermitian time-modulated three-qubit system
Figure 3: Schematic representation of a three-qubit spin system with pairwise $ZZ$ interactions, subject to a transverse magnetic field (simplified here as a perpendicular vector to the plane). The system is governed by the effective NHH in Eq. (1). The magnetic field strength acting on each spin qubit generally varies, as indicated by the differently colored clouds surrounding the qubits. By appropriately time-modulating magnetic fields and qubit interactions, according to Eq. (IV.1), one can realize a nontrivial permutation group acting on the system’s NHH eigenspace.
We now proceed to describe the implementation of a nontrivial state permutation group in the eigenspace of a time-modulated three-qubit system as an illustrative example (see also Fig. 3).
@@ IV.1 Description of the state transfer protocol
Figure 4: Dynamically encircling EPs in a three-qubit system governed by the NHH in Eq. (1) by time modulating the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$, according to Eq. (IV.1), over a winding period $T$. (a) A schematic representation of the initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold. (b),(c) The evolution of various initialized states at times $T/2<t<T$ and at $t=T$, respectively. The Riemann energy sheet configuration shown in panel (b) corresponds to $t=T/2$. (d) Fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and the initial eigenstate $|\psi_{k}\rangle$, $k=1,\dots,8$, of the NHH. System parameters are set as $f_{3}=2f_{1}=2f_{2}=2$, $J_{12}=J_{13}=J_{23}=1$. The time period is $T=1500$ [arb. units], the encircling radius $r=1$, and the initial phase $\phi_{0}=\pi$.
@@
Figure 5: Dynamically spiraling EPs in a three-qubit system governed by the NHH in Eq. (1) by time modulating both the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$ and qubit coupling $J_{13}$, according to Eq. (IV.1), over a winding period $T$. (a) A schematic representation of the initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold. (b),(c) The evolution of various initialized states at times $T/2<t<T$ and at $t=T$, respectively. The Riemann energy sheet configuration shown in panel (b) corresponds to $t=T/2$, which is different from that in panel (a), and which is induced by the additional modulation of the qubit coupling $J_{13}(t)$. The latter enables to couple different Riemann energy sheets compared to that in Fig. 3 with static interactions. (d) Fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and the initial eigenstate $|\psi_{k}\rangle$, ($k=1,\dots,8$), of the given three-qubit NHH. The rest of the system parameters are the same as in Fig. 3. We note that the same result is obtained when modulating the qubit coupling $J_{23}$ in Eq. (IV.1).
@@
In Fig. 2(b) we present a typical energy spectrum for the interacting three qubits, governed by the NHH in Eq. (1). As can be seen from the figure, the winding around EPs in the $(x,y)$ plane alone can only result in the state transfer between eigenstates belonging to individual Riemann pairs. In order to achieve the state flip between different Riemann pairs as well, one has to ensure that these sheets can be ‘transported’ along the $z$-axis in Fig. 2(b). The intended approach is reminiscent to the case studied in Ref. [41], where individual pairs of Riemann manifolds of a bosonic multimode spectrum were manipulated in a similar manner. Here, such control can be achieved by further modulating the interaction terms $J_{mn}$ in the NHH. Specifically, we require the system parameters to be time-modulated as follows:
$x(t)$ $=$ $r\sin(\omega t+\phi_{0}),$
$y(t)$ $=$ $1+r\cos(\omega t+\phi_{0}),$
$J_{mn}(t)$ $=$ $J_{mn}\cos^{2}({\omega t}/{2}),$ (11)
@@
where $r$ is the circle radius in the $(x,y)$ plane, $\omega=2\pi/T$ is an angular frequency determined by a winding period $T$, and $\phi_{0}$ is the initial phase. Moreover, the qubit coupling strengths $J_{mn}$, with $m,n=1,2,3$, and $m<n$, can be modulated individually, i.e., not necessarily simultaneously. To track a state evolution in such a time-modulated system, we numerically solve the Schrödinger equation $i\partial_{t}|\psi(t)\rangle=H|\psi(t)\rangle$ with the Runge-Kutta method.
@@
Alternatively, one could describe the trajectory in Eq. (IV.1) as dynamical spiraling around EPs rather than mere dynamical encirclement. This is because, in addition to orbiting a circular path in the 2D $(x,y)$ parameter space, the system is simultaneously driven along a perpendicular axis, which is defined by the couplings $J_{mn}$, as illustrated in Fig. 2.
@@
IV.2 Dynamical encirclement and spiraling around exceptional points in a three-qubit system
@@
To elucidate the state transfer characterized by the overlap of distinct Riemann pairs during the dynamical encirclement and spiraling of EPs, we give illustrative examples in Figs. 4 and 5, respectively.
@@
When plotting Fig. 4, it is assumed that time modulation is only performed over the parameters $(x,y)$, leaving interactions unchanged $J_{12}=J_{13}=J_{23}=1$, i.e., we solely implement the dynamical encirclement of EPs.
We initialize the systems in one of the NHH eigenstates, as depicted by colored spheres in Fig. 4(a). By starting orbiting around EPs in the $(x,y)$ plane, at the time $t=T/2$ the evolving states cross corresponding energy branch cuts resulting in the state transfer, as shown in panels (b) and (c) in Fig. 4. We calculate the fidelity between the evolving state $\psi(t)$ and initial eigenstates of the NHH at $t=0$ in Fig. 4(d) when EPs encirclement is realized in the counterclockwise direction. We note that identical results are obtained when winding in the opposite, i.e., clockwise, direction, because the non-adiabatic jumps are eliminated in the system due to the real spectrum of the NHH. The resulting permutation can be symbolically written in a cyclic notation as $p=(1,5)(2,6)(3,4)(7,8)$, where the numbers in parentheses account for the cyclic permutations between corresponding eigenvectors. For instance, the notation $(1,5)$ means that the pair of eigenstates $\psi_{1}$ and $\psi_{5}$ are permuted with each other.
@@
It is evident that with the fixed qubit couplings described above, the system is significantly limited in the number of realizable eigenstate flips. To overcome this restriction one may additionally modulate the coupling between, e.g., the first and third qubits, i.e., by modulating the parameter $J_{13}(t)$ in Eq. (IV.1).
@@
We demonstrate the outcome of such a dynamical spiraling around EPs in Fig. 5. The set of possible initialized eigenstates in Fig. 5(a) are the same as in Fig. 4(a). By modulating all three parameters, $[x(t),y(t),J_{13}(t)]$, an alternative combination of state flips can be achieved. Indeed, as illustrated in Fig. 5(b), modulating $J_{13}$ enables the coupling of different pairs of the Riemann manifolds at time $t=T/2$, compared to the configuration shown in Fig. 4(b). Consequently, the following permutation $p=(1,2)(3,6)(4,5)(7,8)$ is implemented, as further confirmed by the calculated state fidelity in Fig. 5(d). We note that the same result can be also achieved when modulating the coupling $J_{23}(t)$ instead.
@@ IV.3 Implementing a nontrivial permutation group in a three-qubit system
By choosing various combinations of the modulated qubit couplings $J_{mn}$, along with modulated parameters $(x,y)$ in Eq. (IV.1), six different permutations can be realized in the given three-qubit system. These permutations can be compactly expressed using cyclic notation as follows:
$p_{1}$ $=$ $(1,5)(2,6)(3,4)(7,8),$
$p_{2}$ $=$ (1,2)(3,6)(4,5)(7,8), if (J13) or $(1,2)(3,6)(4,5)(7,8),\text{ if $(J_{13})$ or $(J_{23})$},$
$p_{3}$ $=$ (1,5)(2,6)(3,7)(4,8), if $(1,5)(2,6)(3,7)(4,8),\text{ if $(J_{13},J_{23})$},$
$p_{4}$ $=$ (1,3)(2,6)(4,5)(7,8), if $(1,3)(2,6)(4,5)(7,8),\text{ if $(J_{12})$},$
$p_{5}$ $=$ (1,3)(2,4)(5,7)(6,8), if (J12,J13) or $(1,3)(2,4)(5,7)(6,8),\text{ if $(J_{12},J_{13})$ or $(J_{12},J_{23})$},$
$p_{6}$ $=$ (1,8)(2,6)(3,7)(4,5), if $(1,8)(2,6)(3,7)(4,5),\text{ if $(J_{12},J_{13},J_{23})$},$ (12)
where the parameters in the parenthesis $(J_{mn},\dots,J_{m^{\prime}n^{\prime}})$ indicate the qubit couplings that must be modulated to realize the corresponding permutation. We numerically confirm these findings in Appendix D.
The generators $p_{i}$ in Eq. (IV.3) form a small non-Abelian permutation group $\cal G$ with 576 elements, and whose group extension can be read as
@@
The core of the group $\cal G$ is formed by the direct product of two alternating groups ${\cal A}_{4}$ of order 12, followed by nested semidirect products of the two cyclic groups ${\cal C}_{2}$ each of order 2.
@@
As such, a non-Abelian permutation group $\cal G$, acting on the eigenspace of the NHH, can be constructed through appropriate time modulations of the system parameters in Eq. (IV.1). This construction utilizes the nontrivial energy topology associated with the EPs of the pseudo-Hermitian Hamiltonian $H$, which effectively suppresses the occurrence of non-adiabatic transitions in the system’s dynamics.

### v2

@@ III.2 Interacting qubits
When the qubit interactions in Eq. (1) are switched on, the NHH spectrum takes the form of a stack of Riemann sheet pairs, determined by the function $\alpha(x,y)$, with the spacing between sheet pairs determined by a function of the $ZZ$ coupling strengths $J_{kl}$, as illustrated in Fig. 2. Hence, for interacting qubits, the EP of a single qubit extends to an exceptional line (EL), as depicted in Fig. 2. This EL is orthogonal to the $(x,y)$ plane and intersects it at the point $(x=0,y=1)$.
To determine whether a non-Hermitian multiqubit system enables a multistate switch, one can examine a topological invariant associated with EPs that is known as chirality [13]. Specifically, when the chirality is non-zero, the eigenstates can be swapped while encircling an EP. In contrast, when the chirality is zero, winding around the EP does not induce a state flip; instead, each eigenstate returns to itself after completing a full encircling cycle [13].
@@ IV Implementing a state permutation group in a non-hermitian time-modulated three-qubit system
Figure 3: Schematic representation of a three-qubit spin system with pairwise ZZZZ interactions, subject to a complex transverse magnetic field (simplified here as a vector perpendicular to the plane). The system is governed by the effective NHH of Eq. (1). In general, the magnetic field strength acting on each spin qubit is different, as indicated by the differently colored clouds surrounding the qubits. By appropriately time-modulating the magnetic fields $\vec{\mathbb{B}}^{k}$ and the qubit interactions $J_{kl}$ of Eq. (1) according to Eq. (IV.1), one can realize a nontrivial permutation group acting on the system’s NHH eigenspace.
We now proceed to describe the implementation of a nontrivial state permutation group in the eigenspace of a time-modulated three-qubit system as an illustrative example. A schematic of the protocol is shown in Fig. 3.
@@ IV.1 Description of the state transfer protocol
Figure 4: Dynamically encircling around EPs in a three-qubit system governed by the NHH in Eq. (1) by time modulating the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$, according to Eq. (IV.1), over a winding period $T$. (a) A schematic representation of the initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold. (b),(c) The evolution of various initialized states at times $T/2<t<T$ and at $t=T$, respectively. (d) Fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and the initial eigenstate $|\psi_{k}\rangle$, $k=1,\dots,8$, of the NHH. System parameters are set as $f_{3}=2f_{1}=2f_{2}=2$, $J_{12}=J_{13}=J_{23}=1$. The time period is $T=2500$ [arb. units], the semi-minor and semi-major axis of the enclosing ellipse are $r_{x}=3$, and $r_{y}=6$, respectively, and the initial phase $\phi_{0}=\pi$.
@@
Figure 5: Dynamically spiraling EPs in a three-qubit system governed by the NHH in Eq. (1) by time modulating both the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$ and qubit coupling $J_{13}$, according to Eq. (IV.1), over a winding period $T$. (a) A schematic representation of the initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold. (b),(c) The evolution of various initialized states at times $T/2<t<T$ and at $t=T$, respectively. The Riemann energy sheet configuration shown in panel (b) corresponds to $t=T/2$ and is distinct from that in panel (a), showing distortions and shifts as a result of the additional modulation of the qubit coupling $J_{13}(t)$. This enables more couplings between different Riemann energy sheets than for the situation in Fig. 3 with static interactions. (d) Fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and the initial eigenstate $|\psi_{k}\rangle$, ($k=1,\dots,8$), of the given three-qubit NHH. The rest of the system parameters are the same as in Fig. 3. We note that the same result is obtained when modulating the qubit coupling $J_{23}$ in Eq. (IV.1).
@@
In Fig. 2(b) we present a typical energy spectrum for three interacting qubits, governed by the NHH of Eq. (1). As can be seen from the figure, winding around EPs that is confined within the $(x,y)$ plane can only result in the state transfer between eigenstates belonging to individual Riemann pairs. In order to achieve state flips between different Riemann pairs as well, one has to ensure that these sheets can be ‘transported’ along the $z$-axis in Fig. 2(b). Our approach is motivated by the programmable switch studied in Ref. [43], where individual pairs of Riemann manifolds of a bosonic multimode spectrum were similarly manipulated. In the current work, such control can be achieved by further modulating the interaction terms $J_{kl}$ in the NHH, Eq. (1). Specifically, we shall require the system parameters to be time-modulated as follows:
$x(t)$ $=$ $r_{x}\sin(\omega t+\phi_{0}),$
$y(t)$ $=$ $r_{y}[1+\cos(\omega t+\phi_{0})],$
$J_{kl}(t)$ $=$ $J_{kl}\cos^{2}({\omega t}/{2}),$ (11)
@@
where $r_{x,y}$ define the semi-minor and/or semi-major axis of the enclosing ellipse in the $(x,y)$ plane, $\omega=2\pi/T$ is an angular frequency determined by a winding period $T$, and $\phi_{0}$ is the initial phase. Clearly, when $r_{x}=r_{y}$, the ellipse loop reduces to a circle. In this work, however, we do not restrict ourselves exclusively to circular orbits. Indeed, the topological nature of the state transfer resulting from dynamically encircling EPs should not depend on the shape of the encircling loop. The qubit coupling strengths $J_{kl}$, with $k,l=1,2,3$, and $k<l$, can be modulated individually, i.e., it is not necessary to modulate them simultaneously. To track the state evolution in such a time-modulated system, we numerically solve the Schrödinger equation $i\partial_{t}|\psi(t)\rangle=H|\psi(t)\rangle$ with the Runge-Kutta method.
@@
Alternatively, one can describe the trajectory in Eq. (IV.1) as dynamical spiraling around EPs rather than mere dynamical encirclement. This is because, in addition to orbiting a circular path in the 2D $(x,y)$ parameter space, the system is simultaneously driven along a perpendicular axis defined by the time-dependent couplings $J_{kl}(t)$, as illustrated in Fig. 2. This dynamical spiraling around the EPs is thus similar to the protocol described in Ref. [43], except for the key fact that in the current situation the dynamics of the system now remain purely adiabatic.
@@
IV.2 Adiabatic dynamical encirclement and spiraling around exceptional points in a three-qubit system
@@
To elucidate the state transfer characterized by the overlap of distinct Riemann pairs during the dynamical encirclement and spiraling of EPs, we now give two illustrative examples for the interacting three-qubit system in Figs. 4 and 5, respectively.
@@
In Fig. 4, we assume that the time modulation is only performed over the parameters $(x,y)$, leaving interactions unchanged $J_{12}=J_{13}=J_{23}=1$, i.e., we implement only the dynamical encirclement of EPs and no spiralling around these.
We initialize the systems in one of the NHH eigenstates, as depicted by the colored spheres in Fig. 4(a). The initial eigenstates $\psi_{i}$ occupy their respective Riemann energy sheets $E_{i}$, which are ordered according to Eq. (7), in the Hermitian phase of the non-Hermitian Hamiltonian $\hat{H}$, where the complex magnetic field $\vec{\mathbb{B}}=0$, i.e., $(x=0,y=0)$. In Figs.  4(a)–(c), the evolving eigenstates $\psi_{i}$ are represented as colored spheres: $\psi_{1}$ (blue), $\psi_{2}$ (green), $\psi_{3}$ (yellow), $\psi_{4}$ (red), $\psi_{5}$ (cyan), $\psi_{6}$ (magenta), $\psi_{7}$ (black), and $\psi_{8}$ (brown).
After initiating the counterclockwise encircling of the EPs in the $(x,y)$ plane, at time $t=T/2$ the evolving states cross the corresponding energy branch cuts, resulting in the state transfer shown in panels (b) and (c) of Fig. 4. Specifically, due to the couplings between energy manifolds $E_{1}\leftrightarrow E_{5}$, $E_{2}\leftrightarrow E_{6}$, $E_{3}\leftrightarrow E_{4}$, and $E_{7}\leftrightarrow E_{8}$ in the region $[x=0,y>1]$, the evolving states transition as follows [see Fig. 4(b)]
$\psi_{1}\rightarrow E_{5},\ \psi_{2}\rightarrow E_{6},\ \psi_{3}\rightarrow E_{4},\ \psi_{4}\rightarrow E_{3},$
$\psi_{5}\rightarrow E_{1},\ \psi_{6}\rightarrow E_{2},\ \psi_{7}\rightarrow E_{8},\ \psi_{8}\rightarrow E_{7}.$ (12)
This obtained configuration of evolving eigenstates remains unchanged for the rest of the evolution until the end of the winding period $t=T$. The resulting permutation can be symbolically written in a cyclic notation as
$p_{1}=(\psi_{1},\psi_{5})(\psi_{2},\psi_{6})(\psi_{3},\psi_{4})(\psi_{7},\psi_{8}),$ (13)
where the numbers in parentheses account for the cyclic permutations between corresponding eigenvectors. For instance, the notation $(\psi_{1},\psi_{5})$ means that the pair of eigenstates $\psi_{1}$ and $\psi_{5}$ are permuted with each other.
Figure 4(d) shows the fidelity between the evolving state $\psi(t)$ and the initial eigenstates of the NHH at $t=0$ when winding around the EPs is realized in the counterclockwise direction. The fidelities show a clear switching behavior between an initial eigenstate at $t=0$ and a different eigenstate at $t=T$, confirming the perfect state transfer between eigenstates. Identical results are obtained when winding in the opposite (i.e., clockwise) direction, because the non-adiabatic jumps, and associated with them chiral mode switching, have been eliminated in the system due to the real spectrum of the NHH.
@@
It is evident that with the fixed qubit couplings described above, while we see perfect eigenstate transfer, the system is still significantly limited in the number of realizable eigenstate flips. To overcome this restriction one can additionally modulate the coupling between, e.g., the first and third qubits, i.e., by modulating the parameter $J_{13}(t)$ in Eq. (IV.1). This gives rise to a dynamical spiraling around the EP, as well as to time-dependent modifications of the Riemannian energy manifolds.
@@
We demonstrate the outcome of such a dynamical spiraling around EPs in Fig. 5. The set of possible initialized eigenstates in Fig. 5(a) are the same as in Fig. 4(a). By modulating all three parameters, $x(t),y(t),J_{13}(t)]$, an alternative combination of state flips can be achieved.
When the system parameters are modulated according to Eq. (IV.1) with the time-dependent coupling $J_{13}(t)$ now included, the Riemann sheets begin to deform and also to shift along the $z$-axis, as shown in Figs. 5(b) and (c). At one half of the winding period, $t=T/2$ [$x=0$, $y>1$], several energy manifolds become coupled [see Fig. 5(b)]:
$E_{1}\leftrightarrow E_{2},\ E_{3}\leftrightarrow E_{6},\ E_{4}\leftrightarrow E_{5},\ E_{7}\leftrightarrow E_{8}.$ (14)
These couplings between energy sheets induce transitions among the eigenstates, analogous to the scenario shown in Fig. 4. However, due to the additional modulation of $J_{13}$, the specific pairings of energy surfaces at $t=T/2$ that are determined by Eq. (14) differ from those in Fig. 4(b). Consequently, via this dynamical spiraling around the EPs, the following permutation is implemented:
$p_{2}=(\psi_{1},\psi_{2})(\psi_{3},\psi_{6})(\psi_{4},\psi_{5})(\psi_{7},\psi_{8}).$ (15)
These state permutations are confirmed by calculations of the state fidelity, depicted in Fig. 5(d). We note furthermore that the same permutation result can also be achieved by modulating the coupling $J_{23}(t)$ instead of $J_{13}(t)$.
@@ IV.3 Implementing a nontrivial permutation group in a three-qubit system
By choosing various combinations of the time-modulated qubit couplings $J_{kl}(t)$, together with time-modulated parameters $(x(t),y(t))$ in Eq. (IV.1), six different permutations can be realized in the given three-qubit system. These permutations can be compactly expressed using cyclic notation as follows:
$p_{1}$ $=$ $(\psi_{1},\psi_{5})(\psi_{2},\psi_{6})(\psi_{3},\psi_{4})(\psi_{7},\psi_{8}),$
$p_{2}$ $=$ $(\psi_{1},\psi_{2})(\psi_{3},\psi_{6})(\psi_{4},\psi_{5})(\psi_{7},\psi_{8}),\text{ if $(J_{13})$ or $(J_{23})$},$
$p_{3}$ $=$ $(\psi_{1},\psi_{3})(\psi_{2},\psi_{6})(\psi_{4},\psi_{8})(\psi_{5},\psi_{7}),\text{ if $(J_{13},J_{23})$},$
$p_{4}$ $=$ $(\psi_{1},\psi_{3})(\psi_{2},\psi_{6})(\psi_{4},\psi_{5})(\psi_{7},\psi_{8}),\text{ if $(J_{12})$},$
$p_{5}$ $=$ $(\psi_{1},\psi_{3})(\psi_{2},\psi_{4})(\psi_{5},\psi_{7})(\psi_{6},\psi_{8}),\text{ if $(J_{12},J_{13})$},$
$p_{6}$ $=$ $(\psi_{1},\psi_{8})(\psi_{2},\psi_{6})(\psi_{3},\psi_{7})(\psi_{4},\psi_{5}),\text{ if $(J_{12},J_{13},J_{23})$},$
where the parameters in the parenthesis $(J_{kl},\dots,J_{k^{\prime}l^{\prime}})$ indicate the qubit couplings that must be modulated to realize the corresponding permutation. We confirm these findings numerically in Appendix C, where we also provide further examples and discussion of the time-dependent distortions of the Riemannian energy surfaces induced by time-modulation of the $J_{kl}(t)$ couplings. We note that the permutation $p_{5}$ in Eq. (IV.3) can be alternatively realized by modulating the qubit couplings $(J_{12},J_{23})$ instead.
The generators $p_{i}$ in Eq. (IV.3) form a small non-Abelian permutation group $\cal G$ with 576 elements, whose group extension can be read as
@@
The core of the group $\cal G$ is formed by the direct product of two alternating groups ${\cal A}_{4}$ of order 12, followed by nested semidirect products of the two cyclic groups ${\cal C}_{2}$ each of order 2 [60].
@@
As such, a non-Abelian permutation group $\cal G$, acting on the eigenspace of the NHH, can be constructed through appropriate time modulations of the system parameters in Eq. (IV.1). This construction utilizes the nontrivial energy topology associated with the EPs of the pseudo-Hermitian Hamiltonian $H$, which effectively suppresses the occurrence of NATs in the system’s dynamics. Notably, these state permutations are topological in nature, since the described state-switching protocol does not depend on the shape of the dynamical loops which enclose the EPs.